OpenStudy (anonymous):

:) I know it's a Saturday but would someone mind explaining this question to me? Bricks are delivered to a work site and stacked in rows and columns, forming a rectangular prism. The length of the prism is 1 foot greater than its width, and its height is 2 feet less than its width. Find the dimensions of the prism formed by the bricks, given that its volume is 40 cubic feet.

OpenStudy (anonymous):

This is what I have so far: l * w * h = V l = w + 1 h = w – 2 (w+1)(w)(w – 2) = 40 ft^3

OpenStudy (anonymous):

Then I have to get 40 on the other side right?

OpenStudy (anonymous):

Then I just get lost...

OpenStudy (***[isuru]***):

first expand the equation u got (w+1)(w)(w – 2) = 40 (w^2 +w)(w - 2) = 40 w^3 - 2w^2 w^2 -2w = 40 w^3 -w^2 -2w - 40 = o now factorize this ( w - 4) (w^2 + 3w + 10 ) = 0 now u have exactly 4 as an 1 answer u can solve the other equation and get another 2 answers if u want hope this will help ya!

OpenStudy (anonymous):

Hold on, is 4 the width? You have to use synthetic division to test possible roots right? And if there's no remainder for a specific number, it's a root?

OpenStudy (anonymous):

so you could make the table to solve for this and anything were the remainder is 0 means it is a factor or root...